A company is currently evaluating a project which requires investments of 5000 now, and 2000 at the end of year. The cash inflows from the project will be 7000 at the end of year 2 and 6000 and the end of year 3. If the discount rate is 16%, what is the net present value of the project?


1
Expert's answer
2021-09-13T00:29:09-0400

"NPV= \sum^n_{i=1} \frac{CF_i}{(1+r)^i} - \sum^n_{i=0} \frac{IC_i}{(1+r)^i} \\
CF_1=0 \\
CF_2=7000 \\
CF_3=6000 \\
IC_0=5000 \\
IC_1=2000 \\
r=16 \% \\
NPV= \frac{7000}{(1+0.16)^2} + \frac{6000}{(1+0.16)^3} - \frac{5000}{1} - \frac{2000}{1+0.16} \\
= 5202.14 + 3843.94 -5000 -1724.13 \\
= 2321.95"